OpenStudy (anonymous):

Find the x- and y intercepts and graph the intercepts. Please show all work. 2x-3y=9

OpenStudy (anonymous):

x=((3y+9)/2)... y=((2x-9)/3)

OpenStudy (anonymous):

x=0 then 2(0)-3y=9 -3y=9 y=9/-3 y=-3 (0,-3)